OpenStudy (anonymous):

Find the zeros of the function. f(x) = 9x^2 + 6x - 8 Please help? This stuff really confuses me

OpenStudy (anonymous):

f(x)=0 9x^2+6x-8=0

OpenStudy (anonymous):

Now, solve for x

OpenStudy (anonymous):

So its 2/3 and -4/3?

OpenStudy (pradipgr817):

\[3x \times(3x+4)-2(3x+4)\] \[(3x-2)(3x+4)=0\] x=2/3,-4/3

OpenStudy (anonymous):

Thank you! Makes much more sense now, im not sure what I was trying to do before

hartnn (hartnn):

yup,both roots are correct :)